OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (-3, 2) and (2, 1). y + 3 = -5(x - 2) y - 2 = -5(x + 3) y + 3 = - (x - 2) y - 2 = - (x + 3)

OpenStudy (ccswims):

first you have to find slope

OpenStudy (anonymous):

Thanks so much @ccswims! You have been extremely helpful helping me understand my more recent questions, and I really appreciate it :)

OpenStudy (ccswims):

awe, no problem!

OpenStudy (anonymous):

Wait, so I'm still a little confused on what answer choice is correct

OpenStudy (ccswims):

ok, I'm just confused on the last 2 options. Can you re-paste or just type in what the last 2 options are again?

OpenStudy (ccswims):

Your slope should be -1/5

OpenStudy (anonymous):

Do you know the slope formula, @lauren24680

OpenStudy (anonymous):

use this formula (y-y1)/(y2-y1)=(x-x1)/(x2-x1)

OpenStudy (anonymous):

where (x1,y1) and (x2,y2) are the given points

OpenStudy (anonymous):

\[m =\frac{ (y2-y2) }{( x2-x1)}\] Find the slope yourself. :)

OpenStudy (anonymous):

you will arrive at the solution directly

OpenStudy (anonymous):

Opps, sorry there was a typo in the last two options @ccswims. There is supposed to be 1/5 behind the - in both of them.

OpenStudy (ccswims):

yea, that's what I though I just wanted to make sure

OpenStudy (ccswims):

So using our rate of change formula we now our slope is -1/5

OpenStudy (ccswims):

So already cross out options A and B

OpenStudy (ccswims):

And D can't be right because point-slope is \[y - y1 = m (x - x1)\] and for the y-axis part they put 2, which is an x-axis

OpenStudy (ccswims):

So what's left?

OpenStudy (anonymous):

(y-y1)/(y2-y1)=(x-x1)/(x2-x1)where (x1,y1) and (x2,y2) are the given points @lauren24680

OpenStudy (ccswims):

^^ yes I showed how to do that in the beginning of the post

OpenStudy (anonymous):

Thank you so much you guys, especially you @ccswims. I understand how to solve it now